OpenStudy (anonymous):

Another practice problems I'm stuck on: logbase3(7x+14)-logbase3(x-3)=2 I can use the property rules to make logbase3((7x+14)/(x-3))=2 And then...??

ganeshie8 (ganeshie8):

good so far, next, change to exponent form

ganeshie8 (ganeshie8):

\(\large \mathbb{\log_a x = b}\) means, \(\large \mathbb{ x = a^b}\)

OpenStudy (shamil98):

\[\large \log_3 \frac{ 7x+14 }{ x-3 } = 2\] let 7x + 14/x-3 = a so using the rule above, what do you get ? \[\large \log_3 a = 2\] (resub the frac back in after you change to exponent form)

OpenStudy (anonymous):



OpenStudy (anonymous):

Oh my goodness, could not decide which was the best response! But that rule logbaseA(x)=b means x=a^b was all I needed to know, and was able to solve it myself. Good practice! Thanks all :)

OpenStudy (anonymous):

answer was 41/2 (or 20.5) btw

OpenStudy (anonymous):

oh and it's already up there... duh!

OpenStudy (anonymous):

yeah ..:)